OpenStudy (anonymous):

hi are there any people can solve this ques. y''+9y=2cos3x+3sinx forthe method of undetermined coefficients

OpenStudy (unklerhaukus):

how you got your complementary solution , i.e. the solution for y''+9y=0 ?

OpenStudy (unklerhaukus):

we have a second order differential equation of this form \[ay''+by'+cy=0\] if we let \[y=e^{mx}\]\[y'=me^{mx}\]\[y''=m^2e^{mx}\] and substitute \[am^2e^{mx}+bme^{mx}+ce^{mx}=0\] \[(am^2+bm+c)e^{mx}=0\] We have a product of two terms If the product of any two terms equals zero , one of the terms must be zero, e^mx is not zero we arrive at \[am^2+bm+c=0\] the auxiliary equation but you dont have to go through this rigmarole every time just jump from \[ay''+by'+cy=0\]\[\qquad\Downarrow\qquad\Downarrow\]to the auxiliary equation \[\qquad\Downarrow\qquad\Downarrow\]\[am^2+bm+c=0\] then you have to find the \(m\)'s